Vector (Cross) Product 1. Find the vector product (2j-2k) x 5k. Sketch all three vectors onto the coordinate system below Answer: 10 Find the vector product of i+4j-3k and -2i+j-5k. Prove that your answer is perpendicular to the first two vectors by using the dot product Answer: -17i+11j+9k or 17i-11j-9k, depending on the order in which you took the cross product. 2.
Answer #1

1)

Vector product of given vectors is

(2] _ 2k) × 5k-10(j × k)-10(k × k) 10(i)-10(0) = 10i

----------------------------------------------

2)

Vector product of given vectors is

{(old{i}+4old{j}-3old{k}) imes(-2old{i}+old{j}-5old{k})=-2(old{i} imesold{i})+(old{i} imesold{j})-5(old{i} imesold{k})-8(old{j} imesold{i})+4(old{j} imesold{j})-20(old{j} imesold{k})+6(old{k} imesold{i})-3(old{k} imesold{j})+15(old{k} imesold{k})}

{(old{i}+4old{j}-3old{k}) imes(-2old{i}+old{j}-5old{k})=-2(0)+(old{k})-5(-old{j})-8(-old{k})+4(0)-20(old{i})+6(old{j})-3(-old{i})+15(0)}

{(old{i}+4old{j}-3old{k}) imes(-2old{i}+old{j}-5old{k})=(old{k})+5(old{j})+8(old{k})-20(old{i})+6(old{j})+3(old{i})

(old{i}+4old{j}-3old{k}) imes(-2old{i}+old{j}-5old{k})=-17old{i}+11old{j}+9old{k}

------------

The resultant vector is 17i + 11j + 9k

Dot products of the resultant vector with the first vector is {(old{i}+4old{j}-3old{k})cdot(-17old{i}+11old{j}+9old{k})=-17(old{i}cdotold{i})+11(old{i}cdotold{j})+9(old{i}cdotold{k})-68(old{j}cdotold{i})+44(old{j}cdotold{j})+36(old{j}cdotold{k})+51(old{k}cdotold{i})+33(old{k}cdotold{j})-27(old{k}cdotold{k})} (i + 4jー3k) . (-17i + 11j 9k)--17(1) ± 11(0) + 9(0) _ 68(0) + 44(1) 36(0) 51(0) 33(0)-27(1)

{(old{i}+4old{j}-3old{k})cdot(-17old{i}+11old{j}+9old{k})=-17+44-27=0

Dot product of the vectors is zero, hence they are perpendicular to each other.

Dot products of the resultant vector with the second vector is {(-2old{i}+old{j}-5old{k})cdot(-17old{i}+11old{j}+9old{k})=34(old{i}cdotold{i})-22(old{i}cdotold{j})-18(old{i}cdotold{k})-17(old{j}cdotold{i})+11(old{j}cdotold{j})+9(old{j}cdotold{k})+85(old{k}cdotold{i})-55(old{k}cdotold{j})-45(old{k}cdotold{k})}

(-2і +j-5k) . (-171+1 1J ± Okl = 34(1) _ 22(0) _ 18(0) _ 17(0) 11 (1) 9(0) + 85(0)-55(0)-15(1)(-2old{i}+old{j}-5old{k})cdot(-17old{i}+11old{j}+9old{k})=34+11-45=0

Dot product of the vectors is zero, hence they are perpendicular to each other.

-------------------------------------------